What is the solution to the system of equations
shown below?
2x- y+3z=8
x-6y-z=0
-6x+3y-9z=24

Answer by richwmiller(17219)

no solution
2x- y+3z=8 times 3
6x-3y+9z=24 try to add to third equation
-6x+3y-9z=24
add
0=48
no solution